OpenStudy (anonymous):

Explain how to get x=3 from the equation: (5x - 1)/2 + (-3x - 1)/4 = 9/2

OpenStudy (saifoo.khan):

\[\frac{5x-1}{2} + \frac{-3x-1}{4} = \frac92\]that?

OpenStudy (anonymous):

Yes

OpenStudy (saifoo.khan):

Ok, find the LCD of the left side. it will be 4. so we can write as, \[\frac{2(5x-1) +(-3x - 1)}{4} = \frac92\]Agree?

OpenStudy (anonymous):

Okay..

OpenStudy (saifoo.khan):

Now try solving further..

OpenStudy (anonymous):

Simplify the top?

OpenStudy (saifoo.khan):

Yes.

OpenStudy (anonymous):

\[(10x-1) +(-3x-1)/4 =9/2\]

OpenStudy (saifoo.khan):

a mistake, it's 10x - 2

OpenStudy (anonymous):

oh yeah, that's what I meant haha

OpenStudy (saifoo.khan):

ok.. move forward..

OpenStudy (anonymous):

\[7x-3/4 = 9/2\]

OpenStudy (saifoo.khan):

good going.. now cross multiply.

OpenStudy (anonymous):